When Katy was 12, her father was three times her age. Now he is twice her age.
How old is Katy now?
my guess:
let x= katy
father=3x

Answer by jim_thompson5910(35256)

"When Katy was 12, her father was three times her age"

So her dad was 12*3 = 36 years old when she was 12.

-------------------------------------------------------

Katy = 12 yrs old
Dad = 36 yrs old

Add on x years

Katy = (12+x) yrs old
Dad = (36+x) yrs old

We want to know when he'll be twice her age, so

(Dad's Age) = 2*(Katy's Age)

36+x = 2*(12+x)

-------------------------------------------------------

Now solve for x


36+x = 2*(12+x)


36+x = 24 + 2x


36 = 24 + 2x-x


36-24 = 2x-x


12 = x


x = 12


In 12 years, Katy will be 12+12 = 24 years old and her dad will be 36+12 = 48 years old (notice how he's twice as old, since 48 = 24*2)